# Complements of discriminants of real parabolic function singularities. II

### v4

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
Figure 9. System of paths for $\bar{f}$ (left) and $f!$ (right)
The virtual function of the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program. Based on the information on the polynomial $\bar{f}$, let us try to find it in this list. According to § 5.2 of [29] there are exactly 1897 virtual functions of class $\Xi_{2}$ with six real critical points. Only 140 of these virtual functions are associated with real polynomials that have a minimum point. Each of these haves, in addition to the minimum, three critical points with Morse index 1 and two points with Morse index 2. Exactly $20=140/7$ of these virtual functions have four negative and two positive critical values at real points. Of these 20 virtual functions, only 12 have all these positive critical values at the critical points with Morse index 2. Thus, the virtual function associated with $\bar{f}$ is among these 12. Upon a special request, our program printed these 12 virtual functions. It turns out that for all of them the intersection index $\langle\Delta_{7},\Delta_{8}\rangle$ of two cycles vanishing at non-real points is equal to 1. In particular, this is true for the real polynomial $\bar{f}$. Using the Lyashko–Looijenga map in the form of [12], (see the proof of Theorem 7 in [29]), we can move this polynomial within the discriminant complement in $\Xi_{2}$ in such a way that all its real critical values stay fixed, but two non-real critical values move along the corresponding paths shown in Fig. 9 (left) almost until the point 0, then travel close to the real axis to the point $1\in{\mathbb{R}}$ and finally meet at this point. Due to the equality $\langle\Delta_{7},\Delta_{8}\rangle=1,$ the critical points corresponding to these values also meet at a real critical point of class $A_{2}$. With a slight further perturbation, we obtain a generic polynomial with eight real critical points, whose six critical values are the same as those $\bar{f}$, and the other two critical points, obtained from the $A_{2}$ point, have real critical values $1-\zeta$ and $1+\zeta$ with some very small $\zeta$. Denote this polynomial by $f!$, see Fig. 9 (right). Its critical point corresponding to the critical value $\varepsilon-1$ and marked with 1 in this figure is a minimum point. The critical points with values marked with 2, 3, and 4 have Morse index 1, and the critical points marked with 7 and 8 have Morse index 2. According to Lemmas V.3.2 and V.3.3 of [23], the newborn critical point with critical value $1-\zeta$ (marked with 5 in the picture) has an odd Morse index, and the point with critical value $1+\zeta$ has an even Morse index. According to Proposition 40 of [29], a generic polynomial of class $\Xi_{2}$ with eight real critical points cannot have more than one point of local extremum, therefore these odd and even Morse indices are in fact equal to one and two, respectively.
@@
The virtual function associated with the polynomial $f!$ is determined by this graph with the above information on the Morse indices of vertices and the numbering of these vertices corresponding to the ascending order of critical points. The vertices with numbers 1–4 correspond to negative critical values, and vertices 5–8 correspond to positive critical values.
According to the previous conditions, the lowest vertex in Fig. 7 has number 1, the vertices in the middle row have numbers 2, 3, 4, and 5, and the vertices in the upper row have the numbers 6, 7, and 8. The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5 of Fig. 9 (right). Indeed, the intersection index $\langle\Delta_{5},\Delta_{6}\rangle$ of cycles vanishing in two newborn Morse critical points is equal to 1, see Lemma V.3.2 in [23]. In particular, the positive critical values of the polynomial $f!$ correspond to all vertices in the top row of Fig. 7 and one vertex of valence three in the middle row.
@@
It remains to prove that the connected component of the discriminant complement containing the polynomial $f!$ is $S(3)$-invariant. Let $L\sim S^{1}$ again be the set of all linear functions ${\mathbb{R}}^{3}\to{\mathbb{R}}$ of unit norm that vanish at the vector $(1,1,1)$.
@@
Proof. The restriction of the polynomial (61) to the plane $\{l=0\}$ is a non-degenerate homogeneous cubic polynomial that vanishes on three real lines, see Fig. 10 (left). The first assertion of the lemma follows from this analogously to the proof of Lemma 39. Let $a$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$. The restriction of the polynomial to the line through 0 and $a$ is a polynomial function of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0. $\Box$
Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin. These numbers $c(l)$ form a lower semi-continuous function on $L$. By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$. The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant. This union is connected and invariant under the action (44) of the group $S(3)$. By Lemma 47, for $l\equiv x-y$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$. The polynomial $F(l)-\delta$ also belongs to this component because all the polynomials F(l)tF(l)-t, where $t\in[\delta,1]$, are non-discriminant. Therefore, the entire component is $S(3)$-invariant. This concludes the proof of Theorem 42 for the virtual component with $\mbox{Card}=1216$.
Components with $\mbox{Card}=1318$ and $844$. Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other critical points with critical value 0. By the versality of the deformation (42), this stratum is a smooth submanifold of codimension four in $\Xi_{2}$. All polynomials $F(l)$ of the form (61) belong to this stratum.
@@
Denote the fiber of this tubular neighborhood over the point $l\in L$ by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
@@
i.e., it can belong to one of the five intervals into which the critical values divide the line $\mathbb{R}$}. For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal. The set of these non-generic polynomials (63) divides the pyramid into six parts. The intersection of the pyramid with a plane perpendicular to its axis looks as the curvilinear triangle shown in Fig. 11 (center). The non-generic polynomials form three segments in it. In particular, the central point of this triangle corresponds to a polynomial $f_{\triangle}$ with equal critical values at all three saddlepoints. The zero-level set of the function $f_{\triangle}-\tau$, where $\tau$ is the critical value, looks as in Fig. 10 (right). These polynomials, representing various sections of the pyramid and corresponding to different values of $\tau$, sweep out its central axis. An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
@@
Consider again the case $l\equiv x-y$. Let $f$ be a point in one of these three domains in the slice $U(l)$ at the point $F(l)$. We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\zeta$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points. Then, acting as in the proof of Lemma 47, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\zeta$ and $\tilde{\omega}_{1}$. For the same reasons as in the proof of Lemma 47, the corresponding critical points meet at a real critical point of class $A_{2}$. Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points. We again obtain a polynomial with Coxeter-Dynkin graph shown in Fig. 7. This polynomial has four critical points with Morse index 1; three of these points arise from the splitting of the $D_{4}^{-}$ singularity. The three corresponding vertices in the middle row of this graph are the leaf vertex and some two of the vertices connected to upper vertices. Only one of these three critical points has a positive critical value. Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex. This choice explicitly determines the virtual component associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement). Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant. It turns out that $\mbox{Card}=844$ when the right-hand vertex in Fig. 7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other cases. $\Box$
The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l).
The union of the contractible domains $V(l)\subset U(l)$ over all $l$ in this circle is a connected set that is $S(3)$-invariant. All of these domains consist of polynomials with $\mbox{Card}=844$. This proves the assertion of Theorem 42 for the virtual components with $\mbox{Card}=844$.
The union of all sets $W(l)$ also is $S(3)$-invariant. It is the space of a fiber bundle over the same circle, with the fiber $W(l)$ consisting of two contractible domains.
@@
The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $D_{4}^{-}$ stratum at the corresponding polynomials $F(l)$.
@@
The singularity of the function $F(l)$ at the origin has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class in some local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$, where $\tilde{l}\approx l$ up to terms of second degree. The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin. The curves on this surface that are distinguished by the conditions
@@
are tangent to the three lines considered in the previous paragraph. Consider the set of points in the four-dimensional slice $U(l)$, which correspond to polynomials, at which the $D_{4}^{-}$ singularity of $F(l)$ splits into a minimum point with negative critical value and three critical points with Morse index 1 and critical value 0. This set is locally homeomorphic to a single interval with an endpoint at the point $\{F(l)\}$: it corresponds to the axis of one of pyramids in Fig. 11 left. Each function $\hat{F}(l)$ from this set has the form $\check{l}^{2}+\varphi(\check{p},\check{q})$ in local coordinates $\check{l},\check{p},\check{q}$, which are very close to $\tilde{l},\tilde{p},\tilde{q}$, where the zero-level set of the function $\varphi$ looks as in Fig. 10 right. The three curves constituting this zero-level set in the surface $\{\check{l}=0\}$ are very close to three curves (65) in the set $\{\tilde{l}=0\}$. In particular, they are in a canonical one-to-one correspondence with these curves, and hence also with three lines of the intersection of the surface (64) and the plane $\ker l$. The three critical points with Morse index 1 occur at their intersections, see Fig. 10 (right). Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at the critical point obtained by this perturbation from the chosen critical point of $\hat{F}(l)$ is positive and the critical values at the other two critical points with Morse index 1 are negative. Thus, there is a canonical one-to-one correspondence between these three domains in $U(l)$ and the choices of some two out of three lines distinguished by the equations (64) and $l=0$.
@@
Proof. All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [20], [12]. The restriction of the polynomial ((66)) to the plane $\ker l$ is a degenerate cubic form; therefore, its singularity at the origin is more complicated than $D_{4}$. According to Corollary 10.4.9 of [22], this singularity is not of class $E_{6}$. Thus, it remains to choose between the classes $D_{5}^{+}$ and $D_{5}^{-}$. In the particular case when
@@
Proof. By the versality property, the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}$. Therefore, it is sufficient to consider only a small five-dimensional transversal slice of this stratum in $\Xi_{2}$ at the point (66). Any such slice realizes a versal deformation of the corresponding singularity. The components of the discriminant complements of simple singularities in the spaces of their versal deformations are listed explicitly in [26]. For $D_{5}^{-}$ singularities, only one of these components consists of functions with $\mbox{Ind}=-3$: it is the component containing the perturbations shown in Fig. 13. According to Theorem 1.7 of [16], this component is contractible. $\Box$
@@
Figure 13. Perturbation of $D_{5}^{-}$ singularity with $\mbox{Ind}=-3$
@@
Proof. We can assume that $l$ is given by (68). Then the set in consideration contains a generic polynomial with the following properties:
@@
  • the critical values at all critical points with Morse index 1 are lower than all critical values at the points with Morse index 2,
@@
Indeed, a polynomial with these properties can be obtained from the perturbation of (66) shown in the middle part of Fig. 13 by a minor additional perturbation that properly orders the critical values. The intersection indices of its vanishing cycles are derived using the Gusein-Zade—A’Campo method [10].
Our program, upon a special request, lists all virtual functions of $P_{8}^{2}$ class satisfying these conditions. It turns out that there are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$

### v5

@@ 4.6. Components of the discriminant complement of the $P_{8}^{2}$ singularities
Figure 9. System of paths for polynomials $\bar{f}$ (left) and $f!$ (right)
The virtual function of the polynomial $\bar{f}$ is among the virtual functions of class $P_{8}^{2}$ listed by our program. Based on the information on the polynomial $\bar{f}$, let us try to find it in this list. According to § 5.2 of [29] there are exactly 1897 virtual functions of class $\Xi_{2}$ with six real critical points. Only 140 of these virtual functions are associated with real polynomials that have a minimum point. Each of these haves, in addition to the minimum, three critical points with Morse index 1 and two points with Morse index 2. Exactly $20=140/7$ of these virtual functions have four negative and two positive critical values at real points. Of these 20 virtual functions, only 12 have all these positive critical values at the critical points with Morse index 2. Thus, the virtual function associated with $\bar{f}$ is among these 12. Upon a special request, our program printed these 12 virtual functions. It turns out that for all of them the intersection index $\langle\Delta_{7},\Delta_{8}\rangle$ of two cycles vanishing at non-real points is equal to 1. In particular, this is true for the real polynomial $\bar{f}$. Using the Lyashko–Looijenga map in the form of [12] (see the proof of Theorem 7 in [29]), we can move this polynomial within the discriminant complement in $\Xi_{2}$ in such a way that all its real critical values stay fixed, but two non-real critical values move along the corresponding paths shown in Fig. 9 (left) almost until the point 0, then travel close to the real axis to the point $1\in{\mathbb{R}}$ and finally meet at this point. Due to the equality $\langle\Delta_{7},\Delta_{8}\rangle=1,$ the critical points corresponding to these values also meet at a real critical point of class $A_{2}$. With a slight further perturbation, we obtain a generic polynomial with eight real critical points, whose six critical values are the same as those $\bar{f}$, and the other two critical points, obtained from the $A_{2}$ point, have real critical values $1-\zeta$ and $1+\zeta$ with some very small $\zeta>0$. Denote this polynomial by $f!$, see Fig. 9 (right). Its critical point corresponding to the critical value $\varepsilon-1$ and marked with 1 in this figure is a minimum point. The critical points with values marked with 2, 3, and 4 have Morse index 1, and the critical points marked with 7 and 8 have Morse index 2. According to Lemmas V.3.2 and V.3.3 of [23], the newborn critical point with critical value $1-\zeta$ (marked with 5 in the picture) has an odd Morse index, and the point with critical value $1+\zeta$ has an even Morse index. According to Proposition 43 of [29], a generic polynomial of class $\Xi_{2}$ with eight real critical points cannot have more than one point of local extremum, therefore these odd and even Morse indices are in fact equal to one and two, respectively.
@@
The virtual function associated with the polynomial $f!$ is determined by this graph supplied with the information on the Morse indices of critical points corresponding to its vertices, and with the numbering of these vertices corresponding to the ascending order of critical values, i.e., the correspondence of these vertices with the black points in Fig. 9 right. The graph and the Morse indices are known to us, while concerning the numbering we know the following. The lowest vertex in Fig. 7 has number 1, the vertices in the middle row have numbers 2, 3, 4, and 5, and the vertices in the upper row have the numbers 6, 7, and 8. The “leaf” vertex of the second row (connected with only one other vertex) cannot correspond to the point 5. Indeed, the intersection index $\langle\Delta_{5},\Delta_{6}\rangle$ of cycles vanishing in two newborn Morse critical points is equal to 1, see Lemma V.3.2 in [23]. In particular, the positive critical values of the polynomial $f!$ correspond to all vertices in the top row of Fig. 7 and one vertex of valence three in the middle row.
@@
It remains to prove that the connected component of the discriminant complement that contains the polynomial $f!$ is $S(3)$-invariant. Let $L\sim S^{1}$ again be the set of all linear functions ${\mathbb{R}}^{3}\to{\mathbb{R}}$ of unit norm that vanish at the vector $(1,1,1)$.
@@
Proof. The restriction of the polynomial (61) to the plane $\{l=0\}$ is a non-degenerate homogeneous cubic polynomial that vanishes on three real lines, see Fig. 10 (left). The first assertion of the lemma follows from this analogously to the proof of Lemma 40. Let $a\in{\mathbb{R}}^{3}$ be an arbitrary critical point of this polynomial outside the plane $\{l=0\}$. The restriction of the polynomial to the line through 0 and $a$ is a polynomial function of degree three with two critical points 0 and $a$, that has a local minimum with critical value 0 at the point 0. $\Box$
The set of all polynomials $F(l)$ is parametrized by the points of the circle $L/{\mathbb{Z}}_{2}\sim S^{1}$, because F(l)=F(l)F(l)=F(-l). Let $c(l)$ be the smallest critical value of the polynomial (61) at its real critical points different from the origin. These numbers $c(l)$ form a lower semi-continuous function on $L/{\mathbb{Z}}_{2}$. By the previous lemma, this function is positive everywhere; therefore, it is separated from zero by a constant $\delta>0$. The union of the points in $\Xi_{2}$ corresponding to the polynomials $F(l)-\delta$ over all $l\in L$ does not intersect the discriminant. This union is connected and invariant under the action (44) of the group $S(3)$. By Lemma 48, for $l\equiv x-y$ the polynomial F(l)1F(l)-1 belongs to the connected component of the discriminant complement with $\mbox{Card}=1216$. The polynomial $F(l)-\delta$ also belongs to this component because all the polynomials F(l)tF(l)-t, where $t\in[\delta,1]$, are non-discriminant. Therefore, the entire component is $S(3)$-invariant. This concludes the proof of Theorem 43 for the virtual component with $\mbox{Card}=1216$.
Components with $\mbox{Card}=1318$ and $844$. Denote by $\{D_{4}^{-}\}$ the stratum of the discriminant in $\Xi_{2}$ that consists of polynomials having a singularity of class $D_{4}^{-}$ and no other real critical points with critical value 0. By the versality of the deformation (42), this stratum is a smooth submanifold of codimension four in $\Xi_{2}$. All polynomials $F(l)$ of the form (61) belong to this stratum.
@@
Denote the fiber of this tubular neighborhood over the point $\{F(l)\}$, $l\in L$, by $U(l)$. For each $l$, define the subset $V(l)\subset U(l)$ as the union of all Morse polynomials $f\in U(l)$ such that
@@
i.e., it can belong to one of the five intervals into which these critical values divide the line $\mathbb{R}$}. For exceptional interior points of the pyramid, the number of intervals is less than five; this occurs when some of critical values of the polynomial are equal. The set of these non-generic polynomials (63) divides the pyramid into six parts. The intersection of the pyramid with a plane perpendicular to its axis looks as the curvilinear triangle shown in Fig. 11 (center). The non-generic polynomials form three segments in it. In particular, the central point of this triangle corresponds to a polynomial $f_{\triangle}$ with equal critical values at all three saddlepoints. The zero-level set of the function $f_{\triangle}-\tau$, where $\tau$ is the critical value, looks as in Fig. 10 (right). These polynomials, representing various sections of the pyramid and corresponding to different values of $\tau$, sweep out its central axis. An arbitrarily small perturbation of the polynomial $f_{\triangle}$ can arrange the critical values of the three saddlepoints in any prescribed order in ${\mathbb{R}}^{1}$.
@@
Consider again the case $l\equiv x-y$. Let $f$ be a point in one of these three domains in the corresponding slice $U(l)$. We assume that it is sufficiently close to $F(l)$, so that the positive critical value $\xi$ at its highest critical point obtained by splitting the $D_{4}^{-}$ singularity is below the smallest value $\tilde{\omega}_{1}$ of all other positive critical values at real critical points. Then, acting as in the proof of Lemma 48, we can move the polynomial $f$ within the discriminant complement in $\Xi_{2}$ in such a way that the critical values at the real critical points remain fixed, and the critical values at the two non-real points converge to a real value between $\xi$ and $\tilde{\omega}_{1}$. For the same reasons as in the proof of Lemma 48, the corresponding critical points meet at a real critical point of class $A_{2}$. Next, we slightly perturb the obtained polynomial so that this point splits into two real critical points. We again obtain a polynomial with Coxeter-Dynkin graph shown in Fig. 7. This polynomial has four critical points with Morse index 1; three of these points arise from the splitting of the $D_{4}^{-}$ singularity. The three corresponding vertices in the middle row of this graph are the leaf vertex and some two of the vertices connected to upper vertices. Only one of these three critical points has a positive critical value. Depending on the domain in the slice containing $f$, the vertex in the graph corresponding to this critical point is or is not the leaf-like right-hand vertex. This choice explicitly determines the virtual component of the virtual function associated with the obtained polynomial (and, consequently, also with the polynomial $f$ that belongs to the same component of the discriminant complement). Starting from an element of this virtual component, our program listed all of its elements, in particular found its Card invariant. It turns out that $\mbox{Card}=844$ when the right-hand vertex in Fig. 7 corresponds to the critical point of $f$ with positive critical value, and $\mbox{Card}=1318$ in the other cases. $\Box$
The union of the contractible domains $V(l)\subset U(l)$ over all $l\in L$ is a connected set that is $S(3)$-invariant. All of these domains consist of polynomials with $\mbox{Card}=844$. This proves the assertion of Theorem 43 for the virtual components with $\mbox{Card}=844$.
The union of all sets $W(l)$ also is $S(3)$-invariant. It is the space of a fiber bundle over the circle $L/{\mathbb{Z}}_{2}$, with the fiber $W(l)$ consisting of two contractible domains.
@@
The monodromy over the cycle $L/{\mathbb{Z}}_{2}\sim S^{1}$ permutes the two components of the sets $W(l)$ in the slices of the $\{D_{4}^{-}\}$ stratum at the corresponding points $\{F(l)\}$.
@@
The singularity of the function $F(l)$ at the origin has the normal form $\tilde{l}^{2}+\tilde{p}^{3}-\tilde{p}\tilde{q}^{2}$ of $D_{4}^{-}$ class in some local coordinates $\tilde{p}$, $\tilde{q}$, $\tilde{l}$, where $\tilde{l}\approx l$ up to terms of second degree. The surface $\{\tilde{l}=0\}$ is tangent to the plane $\{l=0\}$ at the origin. The set $\{F(l)=0\}$ in this surface consists of three curves that are distinguished by the conditions
@@
and are tangent to the three lines considered in the previous paragraph. Consider the set of points in the four-dimensional slice $U(l)$, which correspond to polynomials, at which the $D_{4}^{-}$ singularity of $F(l)$ splits into a minimum point with negative critical value and three critical points with Morse index 1 and critical value 0. This set is locally homeomorphic to a single interval with an endpoint at the point $\{F(l)\}$: it corresponds to the axis of one of pyramids in Fig. 11 left. Each function $\hat{F}(l)$ from this set has the form $\check{l}^{2}+\varphi(\check{p},\check{q})$ in local coordinates $\check{l},\check{p},\check{q}$, which are very close to $\tilde{l},\tilde{p},\tilde{q}$, where the zero-level set of the function $\varphi$ looks as in Fig. 10 right. The three curves constituting this zero-level set in the surface $\{\check{l}=0\}$ are very close to three curves (65) in the set $\{\tilde{l}=0\}$. In particular, they are in a canonical one-to-one correspondence with these curves, and hence also with three lines of the intersection of the surface (64) and the plane $\ker l$. The three critical points with Morse index 1 occur at their intersections, see Fig. 10 (right). Any choice of one of three critical points of $\hat{F}(l)$ with Morse index 1 specifies the domain of the discriminant complement in $U(l)$ that contains the small perturbations of $\hat{F}(l)$ for which the critical value at the critical point obtained by this perturbation from the chosen critical point of $\hat{F}(l)$ is positive and the critical values at the other two critical points with Morse index 1 are negative. Thus, there is a canonical one-to-one correspondence between these three domains in $U(l)$ and the choices of some two out of three lines distinguished by the equations (64) and $l=0$.
@@
Proof. All singularities of corank 2 that can occur in the deformation (42) of $P_{8}$ singularities are of class $D_{4}$, $D_{5}$ and $E_{6}$, see [20], [12]. The restriction of the polynomial ((66)) to the plane $\ker l$ is a degenerate cubic form. Therefore, its singularity at the origin is more complicated than $D_{4}$. According to Corollary 10.4.9 of [22], this singularity is not of class $E_{6}$. Thus, it remains to choose between the classes $D_{5}^{+}$ and $D_{5}^{-}$. In the particular case when
@@
Proof. By the versality property, the discriminant is equisingular along its smooth stratum $\{D_{5}^{-}\}\subset\Xi_{2}$. Therefore, it is sufficient to consider only a small five-dimensional transversal slice of this stratum in $\Xi_{2}$ at the point (66). Any such slice realizes a versal deformation of the corresponding singularity. The components of the discriminant complements of simple singularities in the spaces of their versal deformations are listed explicitly in [26]. For $D_{5}^{-}$ singularities, only one of these components consists of functions with $\mbox{Ind}=-3$: it is the component containing the perturbations shown in Fig. 13. According to Theorem 1.7 of [16], this component is contractible. $\Box$
@@
Figure 13. Perturbation of a $D_{5}^{-}$ singularity with $\mbox{Ind}=-3$
@@
Proposition 55.
Let the linear function $l$ be given by ((68)). Then the set considered Lemmas 53, 54 contains a generic polynomial with the following properties:
@@
  • the critical values at all critical points with Morse index 1 are lower than all critical values at the critical points with Morse index 2,
@@
Proof. A polynomial with these properties can be obtained from the perturbation of the polynomial (66) that is shown in the middle part of Fig. 13 by a minor additional perturbation that properly orders the critical values. The intersection indices of its vanishing cycles are derived using the Gusein-Zade—A’Campo method [10]. $\Box$
Proof of Lemma 54. All these sets corresponding to different $l$ belong to the same connected component of the discriminant complement. Therefore, we can assume that $l$ is given by (68) and hence contains a polynomial satisfying the conditions listed in Proposition 55. Our program, upon a special request, lists all virtual functions of $P_{8}^{2}$ class satisfying these conditions. It turns out that there are exactly six of them, and all six belong to the virtual component with $\mbox{Card}=156.$ $\Box$